# Radial and vertical constraints on the icy origin of HCO in the HD 163296 Protoplanetary Disk

### v1

@@ 5 Conclusions
We have presented the analysis of multiple spectrally and spatially resolved transitions for the para and ortho isomers of HCO toward the Herbig Ae disk around HD 163296. The main conclusions can be summarized as follows:
@@
According to LTE predictions, we determine the column density of HCO as a function of disk radius and, for the first time, we resolve the excitation temperature and the orto-to-para ratio of HCO in HD 163296.
@@
The column density values range from to cm, showing a peak in the inner disk and a plateau in the outer disk, likely related to the positions of the CO snowline and the dust millimeter edge, respectively.
@@
The temperature profile was complemented with empirical measurements of the vertical height of HCO emission. By comparing with predictions of the temperature structure in HD 163296, we find that HCO is tracing a cold region near the midplane ().
@@
An ortho-to-para ratio close to 2.7 remains relatively constant across the disk, which is still consistent with the statistical value of 3.0, considering the uncertainties. However, the OPR profile presents radial modulations, where the lowest OPR values are coincident with HCO emission rings and lower vertical heights.
@@
Constraints from the excitation conditions and the ortho-to-para ratio are consistent with the formation of HCO by both gas-phase and by grain-surface chemistry, especially in the outer disk, where there is an apparent increase in the UV penetration that would favor the photodesorption of HCO from dust grains, but also the formation of radicals due to CO photodissociation.
@@
Our results suggest that the formation of HCO in relatively cold Herbig Ae/Be disks (e.g., HD 163296 and MWC 480) is different than in their warmer analogs characterized by dust cavities and CHOH detections (e.g., IRS 48, HD 100546, and HD 169142). In contrast, the morphology of and the disk-averaged of HCO in HD 163296 shares similarities with results in T Tauri disks.
@@
Regarding the OPR, the only source where a comparison can be made is TW Hya, a T Tauri disk where HCO is thought to be predominantly formed by gas-phase chemistry. Although the OPR values are consistent with in the inner disk of both sources, TW Hya exhibits a monotonic decrease at the outer radius that is not seen for HD 163296, suggesting differences in the chemistry. Dissimilarities in the column density and excitation temperature profiles also support the above. Further multiple-line observations toward more T Tauri and Herbig Ae/Be sources are needed to disentangle the dominant formation pathway of HCO in protoplanetary disks.
@@ Acknowledgments
C.H.-V. acknowledges support from the National Agency for Research and Development (ANID) – Scholarship Program through the Doctorado Nacional grant no. 2021-21212409. V.V.G gratefully acknowledges support from FONDECYT Regular 1221352, and ANID CATA-BASAL project FB210003. L.I.C. acknowledges support from NASA ATP 80NSSC20K0529. L.I.C. also acknowledges support from NSF grant no. AST-2205698, the David and Lucille Packard Foundation, and the Research Corporation for Scientific Advancement Cottrell Scholar Award.
@@ Appendix A Observational details
HCO Line Robust UV taper JvM () ALMA band 3 o-HCO ALMA band 6 p-HCO o-HCO ALMA band 7 p-HCO p-HCO p-HCO o-HCO
@@ Appendix B Voigt profile line modeling
Following the analysis described in Section 3.4, we model the disk-averaged emission using Voigt profiles instead of Gaussian to better characterize the line wings. The retrieved disk-averaged spectra and the best-fit parameters are shown in Figure 8, where the best-fit models are overlayed with different colors for each HCO isomer. A similar approach was applied in TW Hya for CO observations likely affected by line pressure broadening (Yoshida et al., 2022), which enabled the estimation of its gas surface density. Therefore, the presence of non-Gaussian profiles in HD 163296 could have some implications in constraining its physical conditions, but this is beyond the scope of this work.
@@ Appendix C Optical depth radial profiles
Figure 9 shows the deprojected radial profiles of the line optical depth for all HCO transitions used in this work, resulting from the analysis described in Section 3.4.2. Each color represents a different transition. The two weakest p-HCO lines ( and ) have very similar values for , hence the two curves are overlapping.
Appendix D HCO formation temperature

### v2

@@ 5 Conclusions
We have presented the analysis of multiple spectrally and spatially resolved transitions for the para and ortho isomers of H2CO toward the Herbig Ae disk around HD 163296. The main conclusions can be summarized as follows:
@@
According to LTE predictions, we determine the column density of H2CO as a function of disk radius, and for the first time, we resolve the excitation temperature and the ortho-to-para ratio of H2CO in HD 163296.
@@
The column density values range from to cm-2, showing a peak in the inner disk and a plateau in the outer disk, likely related to the positions of the CO snowline and the dust millimeter edge, respectively.
@@
The temperature profile was complemented with empirical measurements of the vertical height of H2CO emission. By comparing with predictions of the temperature structure in HD 163296, we find that H2CO is tracing a cold region near the midplane ().
@@
An ortho-to-para ratio close to 2.7 remains relatively constant across the disk, which is still consistent with the statistical value of 3.0, considering the uncertainties. However, the OPR profile presents radial modulations, where the lowest OPR values are coincident with H2CO emission rings and lower vertical heights.
@@
Constraints from the excitation conditions and the ortho-to-para ratio are consistent with the formation of H2CO by both gas-phase and by grain-surface chemistry, especially in the outer disk, where there is an apparent increase in the UV penetration that would favor the photodesorption of H2CO from dust grains, but also the formation of radicals due to CO photodissociation.
@@
Our results suggest that the formation of H2CO in relatively cold Herbig Ae/Be disks (e.g., HD 163296 and MWC 480) is different from that in their warmer analogs characterized by dust cavities and CH3OH detections (e.g., IRS 48, HD 100546, and HD 169142). In contrast, the morphology of and the disk-averaged of H2CO in HD 163296 shares similarities with the results in T Tauri disks.
@@
Regarding the OPR, the only source where a comparison can be made is TW Hya, a T Tauri disk where H2CO is thought to be predominantly formed by gas-phase chemistry. Although the OPR values are consistent with in the inner disk of both sources, TW Hya exhibits a monotonic decrease at the outer radius that is not seen for HD 163296, suggesting differences in the chemistry. Dissimilarities in the column density and excitation temperature profiles also support the above. Further multiple-line observations toward more T Tauri and Herbig Ae/Be sources are needed to disentangle the dominant formation pathway of H2CO in protoplanetary disks.
@@ Acknowledgments
We thank the anonymous reviewer for providing useful comments and suggestions. C.H.-V. acknowledges support from the National Agency for Research and Development (ANID) – Scholarship Program through the Doctorado Nacional grant No. 2021-21212409. V.V.G gratefully acknowledges support from FONDECYT Regular 1221352, and ANID CATA-BASAL project FB210003. L.I.C. acknowledges support from NASA ATP 80NSSC20K0529. L.I.C. also acknowledges support from NSF grant no. AST-2205698, the David and Lucille Packard Foundation, and the Research Corporation for Scientific Advancement Cottrell Scholar Award.
@@ Appendix A Observational details
Table 4 summarizes the imaging parameters to obtain a common circularized beam of for all transitions used in this work and the parameter to account for the JvM effect.
H2CO Line Robust UV taper JvM (′′) ALMA band 3 o-H2CO ALMA band 6 p-H2CO o-H2CO ALMA band 7 p-H2CO p-H2CO p-H2CO o-H2CO
@@ Appendix B Voigt profile line modeling
Following the analysis described in Section 3.4, we model the disk-averaged emission using Voigt profiles instead of Gaussian to better characterize the line wings. The retrieved disk-averaged spectra and the best-fit parameters are shown in Figure 8, where the best-fit models are overlaid with different colors for each H2CO isomer. A similar approach was applied in TW Hya for 12CO observations likely affected by line pressure broadening (Yoshida et al., 2022), which enabled the estimation of its gas surface density. Therefore, the presence of non-Gaussian profiles in HD 163296 could have some implications in constraining its physical conditions, but this is beyond the scope of this work.
@@ Appendix C Optical depth radial profiles
Figure 9 shows the deprojected radial profiles of the line optical depth for all H2CO transitions used in this work, resulting from the analysis described in Section 3.4.2. Each color represents a different transition. The two weakest p-H2CO lines ( and ) have very similar values for ; hence, the two curves are overlapping.
Appendix D H2CO formation temperature